OpenStudy (anonymous):

What is the slope of a line that passes through the point (−1, 1) and is parallel to a line that passes through (4, 6) and (−1, −4)?

OpenStudy (dls):

Parallel lines share the same slope so just find the slope of the line passing through (4,6) and (-1,-4) using the old slope formula i.e. \[\Huge Slope(m)=\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (dls):

x1=4 x2=-1 y1=6 y2=-4

OpenStudy (mathmale):

Find this slope (of the given line). Next, determine the slope of a line perpendicular to the first line. How would you do that?

OpenStudy (anonymous):

ok wait a second i'm working it out

OpenStudy (anonymous):

@DLS Is the slope m=-10/5?

OpenStudy (dls):

no 10/5=2

OpenStudy (anonymous):

how did you get that? I followed the equation

OpenStudy (dls):

Nope you made a mistake.\[\Huge \frac{-4-6}{-1-4 } = \frac{\cancel{-}10}{\cancel{-}5}=2\]

OpenStudy (anonymous):

so the slope is 2?

OpenStudy (dls):

yes